# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix D Proof of Lipschitz-continuity Bound
Definition D.2 (Total Variation Distance).
@@
Theorem D.3 ($\widehat{Q}_{k}^{T}$ is $\frac{2}{1-\gamma}\|r_{l}(\cdot,\cdot)\|_{\infty}$-Lipschitz continuous with respect to $F_{z_{\Delta}}$ in total variation distance).
@@
$\left|\widehat{Q}^{T}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}^{T}_{k^{\prime}}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|\leq\left(\sum_{t=0}^{T-1}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
@@
We prove this inductively. First, note that $\widehat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\widehat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$ from the initialization step, which proves the lemma for $T=0$, since $\mathrm{TV}(\cdot,\cdot)\geq 0$. At $T=1$:
$|\widehat{Q}_{k}^{1}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}_{k^{\prime}}^{1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$ $=\left|\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{0}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{\mathcal{T}}_{k^{\prime}}\widehat{Q}_{k^{\prime}}^{0}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|$
$=\bigg{|}r(s_{g},F_{s_{\Delta}},a_{g})+\gamma\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g}}\widehat{Q}_{k}^{0}(s_{g}^{\prime},F_{s^{\prime}_{\Delta}},a^{\prime}_{g})$
$\quad\quad\quad\quad-r(s_{g},F_{s_{\Delta^{\prime}}},a_{g})-\gamma\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g}}\widehat{Q}_{k^{\prime}}^{0}(s^{\prime}_{g},F_{s^{\prime}_{\Delta^{\prime}}},a^{\prime}_{g})\bigg{|}$
@@
In the first and second equalities, we use the time evolution property of $\widehat{Q}_{k}^{1}$ and $\widehat{Q}_{k^{\prime}}^{1}$ by applying the adapted Bellman operators $\widehat{\mathcal{T}}_{k}$ and $\widehat{\mathcal{T}}_{k^{\prime}}$ to $\widehat{Q}_{k}^{0}$ and $\widehat{Q}_{k^{\prime}}^{0}$, respectively, and expanding. In the third and fourth equalities, we note that $\widehat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\widehat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$, and subtract the common ‘global component’ of the reward function.
@@
$|\widehat{Q}^{1}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}_{k^{\prime}}^{1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$ $=\left|\mathbb{E}_{z_{l}\sim F_{z_{\Delta}}}\left[\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\mathbbm{1}\{z_{l}=z\}\right]-\mathbb{E}_{z_{l}^{\prime}\sim F_{z_{\Delta^{\prime}}}}\left[\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\mathbbm{1}\{z_{l}^{\prime}=z\}\right]\right|$
@@
The second equality follows from the linearity of expectations, and the third equality follows by noting that for any random variable $X\sim\mathcal{X}$, $\mathbb{E}_{X}\mathbbm{1}[X=x]=\Pr[X=x]$. The first inequality follows from an application of the triangle inequality and the Cauchy-Schwarz inequality, and the second inequality uses the definition of TV distance. Thus, at $T=1$, $\widehat{Q}$ is $(2\|r_{l}(\cdot,\cdot)\|_{\infty})$-Lipschitz continuous in TV distance, proving the base case.
@@
$\left|\widehat{Q}^{T}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}_{k^{\prime}}^{T}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|\leq\left(\sum_{t=0}^{T-1}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
@@
$|\widehat{Q}^{T+1}_{k}(s_{g},a_{g},F_{z_{\Delta}})$ $-\widehat{Q}_{k^{\prime}}^{T+1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$
$\leq\bigg{|}\frac{1}{|\Delta|}\sum_{i\in\Delta}r_{l}(s_{g},z_{i})-\frac{1}{|\Delta^{\prime}|}\sum_{i\in\Delta^{\prime}}r_{l}(s_{g},z_{i})\bigg{|}$
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta}^{\prime}})-\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta^{\prime}}^{\prime}})\bigg{|}$
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s^{\prime}_{\Delta})\sim\mathcal{J}_{k}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta}^{\prime}})-\mathbb{E}_{(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}})\sim\mathcal{J}_{k^{\prime}}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta^{\prime}}^{\prime}})\bigg{|}$
$\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})+\gamma\left(\sum_{\tau=0}^{T-1}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
$=\left(\sum_{\tau=0}^{T}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
In the first inequality, we rewrite the expectations over the states as the expectation over the joint transition probabilities. The second inequality then follows from Lemma D.12. To apply it to Lemma D.12, we conflate the joint expectation over $(s_{g},s_{\Delta\cup\Delta^{\prime}})$ and reduce it back to the original form of its expectation. Finally, the third inequality follows from Lemma D.5.
@@
Lemma D.5.
For all $T\in\mathbb{N}$, for any $a_{g}\in\mathcal{A}_{g},s_{g}\in\mathcal{S}_{g},s_{\Delta}\in\mathcal{S}_{l}^{k},a_{\Delta}\in\mathcal{A}_{l}^{k},a^{\prime}_{\Delta}\in\mathcal{A}_{l}^{k}$, and for all joint stochastic kernels $\mathcal{J}_{k}$ as defined in Definition D.4:
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{T}(s^{\prime}_{g},a_{g}^{\prime},F_{z^{\prime}_{\Delta}})$ $-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a_{\Delta^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime}_{g},a_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}})\bigg{|}$
$\quad\quad\quad\quad\quad\leq\left(\sum_{\tau=0}^{T-1}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
@@
We prove this inductively. At $T=0$, the statement is true since $\widehat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\widehat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$ and $\mathrm{TV}(\cdot,\cdot)\geq 0$. At $T=1$,
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{1}(s^{\prime}_{g},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\widehat{Q}_{k^{\prime}}^{1}(s^{\prime}_{g},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})\bigg{|}$
@@
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\!\max_{a_{\Delta}^{\prime}}\frac{\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k}-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a^{\prime}_{\Delta^{\prime}}}\frac{\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k^{\prime}}\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
@@
$\max_{a_{\Delta}^{\prime}}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$ $=\sum_{i\in\Delta}\max_{a_{\Delta}^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$
$=\sum_{i\in\Delta}\max_{a^{\prime}_{i}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$
$=\sum_{i\in\Delta}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})$
where $\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime}):=\max_{a_{i}^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$.
Then, analogously to an argument in Lemma B.3 of Anand & Qu (2024), we have:
@@
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
@@
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{x\in\mathcal{S}_{l}}\widetilde{r}_{l}(x,s_{g}^{\prime})F_{s^{\prime}_{\Delta}}(x)-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{x\in\mathcal{S}_{l}}\widetilde{r}_{l}(x,s_{g}^{\prime})F_{s^{\prime}_{\Delta^{\prime}}}(x)\bigg{|}$

### v3

@@ Appendix E Proof of Lipschitz-continuity Bound
Definition E.2 (Total Variation Distance).
@@
Theorem E.3 ($\hat{Q}_{k}^{T}$ is $\frac{2}{1-\gamma}\|r_{l}(\cdot,\cdot)\|_{\infty}$-Lipschitz continuous with respect to $F_{z_{\Delta}}$ in total variation distance).
@@
$\left|\hat{Q}^{T}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\hat{Q}^{T}_{k^{\prime}}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|\leq\left(\sum_{t=0}^{T-1}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
@@
We prove this inductively. First, note that $\hat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\hat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$ from the initialization step, which proves the lemma for $T=0$, since $\mathrm{TV}(\cdot,\cdot)\geq 0$. At $T=1$:
$|\hat{Q}_{k}^{1}(s_{g},a_{g},F_{z_{\Delta}})-\hat{Q}_{k^{\prime}}^{1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$ $=\left|\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{0}(s_{g},a_{g},F_{z_{\Delta}})-\hat{\mathcal{T}}_{k^{\prime}}\hat{Q}_{k^{\prime}}^{0}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|$
$=\bigg{|}r(s_{g},F_{s_{\Delta}},a_{g})+\gamma\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g}}\hat{Q}_{k}^{0}(s_{g}^{\prime},F_{s^{\prime}_{\Delta}},a^{\prime}_{g})$
$\quad\quad-r(s_{g},F_{s_{\Delta^{\prime}}},a_{g})-\gamma\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g}}\hat{Q}_{k^{\prime}}^{0}(s^{\prime}_{g},F_{s^{\prime}_{\Delta^{\prime}}},a^{\prime}_{g})\bigg{|}$
@@
In the first and second equalities, we use the time evolution property of $\hat{Q}_{k}^{1}$ and $\hat{Q}_{k^{\prime}}^{1}$ by applying the adapted Bellman operators $\hat{\mathcal{T}}_{k}$ and $\hat{\mathcal{T}}_{k^{\prime}}$ to $\hat{Q}_{k}^{0}$ and $\hat{Q}_{k^{\prime}}^{0}$, respectively, and expanding. In the third and fourth equalities, we note that $\hat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\hat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$, and subtract the common ‘global component’ of the reward function.
@@
$|\hat{Q}^{1}_{k}(s_{g},a_{g},F_{z_{\Delta}})$ $-\hat{Q}_{k^{\prime}}^{1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$
$=\left|\mathbb{E}_{z_{l}\sim F_{z_{\Delta}}}\left[\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\mathbbm{1}\{z_{l}=z\}\right]-\mathbb{E}_{z_{l}^{\prime}\sim F_{z_{\Delta^{\prime}}}}\left[\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\mathbbm{1}\{z_{l}^{\prime}=z\}\right]\right|$
@@
The second equality follows from the linearity of expectations, and the third equality follows by noting that for any random variable $X\sim\mathcal{X}$, $\mathbb{E}_{X}\mathbbm{1}[X=x]=\Pr[X=x]$. The first inequality follows from an application of the triangle inequality and the Cauchy-Schwarz inequality, and the second inequality uses the definition of TV distance. Thus, at $T=1$, $\hat{Q}$ is $(2\|r_{l}(\cdot,\cdot)\|_{\infty})$-Lipschitz continuous in TV distance, proving the base case.
@@
$\left|\hat{Q}^{T}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\hat{Q}_{k^{\prime}}^{T}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|\leq\left(\sum_{t=0}^{T-1}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
@@
$|\hat{Q}^{T+1}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\hat{Q}_{k^{\prime}}^{T+1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$
$\quad\leq\bigg{|}\frac{1}{|\Delta|}\sum_{i\in\Delta}r_{l}(s_{g},z_{i})-\frac{1}{|\Delta^{\prime}|}\sum_{i\in\Delta^{\prime}}r_{l}(s_{g},z_{i})\bigg{|}$
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta}^{\prime}})-\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta^{\prime}}^{\prime}})\bigg{|}$
$\quad\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s^{\prime}_{\Delta})\sim\mathcal{J}_{k}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta}^{\prime}})-\mathbb{E}_{(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}})\sim\mathcal{J}_{k^{\prime}}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta^{\prime}}^{\prime}})\bigg{|}$
$\quad\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})+\gamma\left(\sum_{\tau=0}^{T-1}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
$\quad=\left(\sum_{\tau=0}^{T}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
In the first inequality, we rewrite the expectations over the states as the expectation over the joint transition probabilities. The second inequality then follows from Lemma E.14. To apply it to Lemma E.14, we conflate the joint expectation over $(s_{g},s_{\Delta\cup\Delta^{\prime}})$ and reduce it back to the original form of its expectation. Finally, the third inequality follows from Lemma E.5.
@@
Lemma E.5.
For all $T\in\mathbb{N}$, for any $a_{g}\in\mathcal{A}_{g},s_{g}\in\mathcal{S}_{g},s_{\Delta}\in\mathcal{S}_{l}^{k},a_{\Delta}\in\mathcal{A}_{l}^{k},a^{\prime}_{\Delta}\in\mathcal{A}_{l}^{k}$, and for all joint stochastic kernels $\mathcal{J}_{k}$ as defined in Definition E.4:
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\hat{Q}_{k}^{T}(s^{\prime}_{g},a_{g}^{\prime},F_{z^{\prime}_{\Delta}})$
$\!-\!\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a_{\Delta^{\prime}}}\hat{Q}_{k^{\prime}}^{T}(s^{\prime}_{g},a_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}})\bigg{|}\!\leq\!\left(\sum_{\tau=0}^{T-1}2\gamma^{\tau}\right)\!\|r_{l}(\cdot,\cdot)\|_{\infty}\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
@@
We prove this inductively. At $T=0$, the statement is true since $\hat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\hat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$ and $\mathrm{TV}(\cdot,\cdot)\geq 0$.
At $T=1$,
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\hat{Q}_{k}^{1}(s^{\prime}_{g},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\hat{Q}_{k^{\prime}}^{1}(s^{\prime}_{g},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})\bigg{|}$
@@
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\!\max_{a_{\Delta}^{\prime}}\frac{\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a^{\prime}_{\Delta^{\prime}}}\frac{\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k^{\prime}}\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}\tilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\tilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
@@
$\max_{a_{\Delta}^{\prime}}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$ $=\sum_{i\in\Delta}\max_{a_{\Delta}^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})=\sum_{i\in\Delta}\max_{a^{\prime}_{i}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})=\sum_{i\in\Delta}\tilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime}),$
where $\tilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime}):=\max_{a_{i}^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$.
Then, we have:
@@
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}\tilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\tilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
@@
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{x\in\mathcal{S}_{l}}\tilde{r}_{l}(x,s_{g}^{\prime})F_{s^{\prime}_{\Delta}}(x)$